How do you solve -6=-(x-5)-3(5+2x)-4(2x-4)? I am totally lost
Answer by josgarithmetic(39617)

Perform the operations and simplify; isolate all terms of x and finish solving for x.

Doing the operations:
-6=-x+5-15-6x-8x+16

Some simplifying:
-6=x-6x-8x+5-15+16
-6=-13x+6
.
.
You finish!